OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 The solution to the system of equations is _____. (13/34, -15/34) (3/22, -7/22) (13/22, 1/22)

OpenStudy (anonymous):

are you asking for someone to check or are those the options?

OpenStudy (anonymous):

those are options

OpenStudy (anonymous):

it is (13/22, 1/22)

OpenStudy (anonymous):

3x - 4y = 7 3x + 2y = -5 When the second equation is subtracted from the first, the resulting equation is _____. -6y= -12 -6y= 12 -2y= -12 2y= 12

OpenStudy (anonymous):

the equations you put are 7x-3y=4 and 2x-4y=1 so if you multiply the first one by 4 and the second one by 3(to get the ys equal) you get 28x-12y=16 and 6x-12y=3 so you would subtract the second one from the first one and get 22x=13 then divide by 22 and get x=13/22.

OpenStudy (anonymous):

i understand that i need help with other one

OpenStudy (anonymous):

oh! okay, well you would set it up the same way then..

OpenStudy (anonymous):

3x-4y=7 3x+2y=-5 you would subtract the second from the first to get rid of x.